OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 2 > -4 and 3x less than or greater to 15. A number line with an open circle on -2, shading to the left, and a closed circle on 5, shading to the right. A number line with a closed circle on -2, shading to the left, and an open circle on 5, shading to the right. A number line with an open circle on -2, a closed circle on 5, and shading in between. A number line with a closed circle on -2, an open circle on 5, and shading in between.

OpenStudy (anonymous):

x - 2 > -4 (add 2 to both sides) x > -4 + 2 x > -2 3x < = 15 ....I am assuming this is less then or equal...is that correct ?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

the less than or equal two sign has this sign < and with a line under it

OpenStudy (anonymous):

the great than or equal two is: > with a line under it

OpenStudy (anonymous):

3x <= 15 (divide by 3) x <= 5 so we have : x > -2 and x <= 5 Learn this... if there is an equal sign, the circle is closed if there is no equal sign, the circle is open < (less then) means shading to the left > (greater then) means shading to the right can you figure the answer from here ?

OpenStudy (anonymous):

,so wich one is it

OpenStudy (anonymous):

Yukisama..I understand what the signs mean, it is in the second inequality that it says " less then or greater " that confused me.

OpenStudy (anonymous):

ah okay :3

OpenStudy (anonymous):

just don't need people confused i like to help out

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

x > -2 ...open circle on -2, shading to the right x <= 5 ..closed circle on 5, shading to the left so your answer is : open circle on -2, closed circle on 5, and shading in between

OpenStudy (anonymous):

can u help with another question

OpenStudy (anonymous):

what you got ?

OpenStudy (anonymous):

Solve for r: C = 2pir. r = C over 2 pi r = 2 c all over pi r = C over pi r = pi times C all over 2

OpenStudy (anonymous):

i think its the second one not sure

OpenStudy (anonymous):

C = 2* pi * r (solve for r) -- divide by 2 * pi C/2*pi = r

OpenStudy (anonymous):

so would tht be the snwer

OpenStudy (anonymous):

no..answer is the first one

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

no problem :)